A 150 foot long footing has 3 ea. #5 reinforcing bars continuous the length of the footing. The bars are 20 foot long each and the design requires a 30 bar diameter lap splice at each splice. If #5 reinforcing weighs 1.04 pounds per linear foot and has a bar diameter of 5/8 inch, how many tons of rebar are in the footing? (Note: 1 ton = 2,000 pounds) Disregard concrete cover.
Answer #1

Hi,

Solution as follows:

  1. Lap Length Calculation:

Bar Diameter = 5/8 inch

Lap Length as per design = 30 time diameter

Lap Length in inches = 5/8 x 30 = 18.75 inch

Lap Length in feet = 18.75/12 = 1.56 feet

  1. Effective length of each bar:

Total Length of one Rebar = 20 ft

Lap Length = 1.56 ft

Effective length = 20 – 1.56 = 18.44 ft

  1. Required Number of Bars and laps:

Hence the required number of bars in single run would be 9. 8 bars will be of full length and one bar will be short length. Accordingly number of laps in single run would be 8.

  1. Rebar Schedule:

Rebar schedule of footings is given below:

TABLE 1: REBAR QUANTITY OF FOOTING

Sl. No.

Particulars

Bar No.

Bar spacing (ft)

Spacing Distance (ft)

Each Bar Length (ft.)

No. of Bars

Unit Wt (lbs per ft.)

Total Weight (lbs)

1

Continuous bar

#5

150.00

3

1.040

468.00

2

Overlap

#5

1.56

24

1.040

38.94

Total weight in pounds

506.94

Since 1 ton = 2000 pounds, rebar weight in ton

0.253

Hope the solution provided above is clear. In case of any clarifications, please leave comment in comments section.

Please give us a good rating, if you are satisfied with the response.

How would you rate our reply?

Keep Chegging

Best Regards

Chegg Expert Team